# Canonical decompositions of 3-connected graphs

### v2

@@ 2.2.2. Splitting stars
Let $(T,{{\mathcal{V}}})$ be a tree-decomposition of $G$, and let $S$ denote the set of induced separations of $(T,{{\mathcal{V}}})$. For every node $t\in T$, let $\sigma_{t}$ denote star of separations induced by the edges of $T$ incident with $t$ and directed to $t$. The splitting stars of $S$ are precisely the stars $\sigma_{t}$ with $t\in T$.
@@
(1$\to$(2). Let $\sigma$ be a splitting star of $N$, and assume for a contradiction that $(A,B)\in N$ interlaces $\sigma$. Then there is $(C,D)\in\sigma$ such that $(A,B)\leqslant(C,D)$ or $(B,A)\leqslant(C,D)$. But we also have $(C,D)<(A,B)$ or $(C,D)<(B,A)$ since $(A,B)$ interlaces $\sigma$. In two cases we obtain immediate contradictions, and in the other two cases we obtain $A\subseteq B$ or $B\subseteq A$ which contradicts the definition of a separation.
(2$\to$(1). Assume that no element of $N$ interlaces $\sigma$; we show that $\sigma$ is a splitting star of $N$. Let $(A,B)\in N$, and assume for a contradiction that there is no $(C,D)\in\sigma$ such that $(A,B)\leqslant(C,D)$ or $(B,A)\leqslant(C,D)$. Then, since $N$ is nested, for every $(C,D)\in\sigma$ we have $(C,D)<(A,B)$ or $(C,D)<(B,A)$, so $(A,B)$ interlaces $\sigma$. ∎
@@
Assume for a contradiction that some mixed-separation $(A,B)$ of $G$ interlaces two distinct splitting stars $\sigma_{1}$ and $\sigma_{2}$ of $M$. By Section 2.2.2, there exist $(C_{1},D_{1})\in\sigma_{1}$ and $(C_{2},D_{2})\in\sigma_{2}$ such that $(D_{1},C_{1})\leqslant(C_{2},D_{2})$. Since $(A,B)$ interlaces $\sigma_{1}$ and $\sigma_{2}$, and since $(A,B)\not<(A,B)$ nor $(B,A)\not<(B,A)$, we either have
@@
2.2.3. Torsos
${}^{+}$ of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$ and no two edges in $E(A\setminus B,B\setminus A)$ share endvertices. We stress that we allow $A\setminus B$ and $B\setminus A$ to be empty. All the usual concepts for mixed-separations extend to ${}^{+}$ in the obvious way.
@@ 2.2.3. Torsos
All nontrivial mixed 3-separations of a 3-connected graph are mixed ${}^{+}$ by Section 1.2. Pairs $(A,V(G))$ for $A\subseteq V(G)$ are ${}^{+}$ but not separations.
Let $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ be a star of ${}^{+}$ of a graph $G$, with leaf-sides $A_{i}$. The bag of $\sigma$ is the intersection $\bigcap_{i\in I}B_{i}$ of all non-leaf sides $B_{i}$. We follow the convention that the bag of the empty star is equal to the vertex-set of $G$.
If all $(A_{i},B_{i})$ are ${}^{+}$, then the torso of $\sigma$ is the graph obtained from $G[\beta]$, where $\beta$ is the bag of $\sigma$, by making every separator $A_{i}\cap B_{i}$ into a clique (by adding all possible edges inside $A_{i}\cap B_{i}$ for all $i\in I$). In general, however, there are (at least) two ways how the notion of a torso can be generalised to stars of ${}^{+}$. Here we present two ways, supported by Figure 2.2.
@@
The expanded torso of $\sigma$ is the graph that is obtained from $G$ as follows. We obtain $(A_{i}^{\prime},B_{i}^{\prime})$ from $(A_{i},B_{i})\in\sigma$ by letting $A_{i}^{\prime}:=A_{i}$ and we obtain $B_{i}^{\prime}$ from $B_{i}$ by adding all endvertices of edges in the separator of $(A_{i},B_{i})$. Then $(A_{i}^{\prime},B_{i}^{\prime})$ is a ${}^{+}$ with the same order as $(A_{i},B_{i})$. We take the torso of the star $\{\,(A_{i}^{\prime},B_{i}^{\prime}):i\in I\,\}$ of ${}^{+}$ as the expanded torso of $\sigma$.
Figure 2.2. Left: the star $\sigma=\{\,(A,B),(C,D)\,\}$. Middle: the expanded torso of $\sigma$. Right: the compressed torso of $\sigma$.
Note that the compressed torso can be obtained from the expanded torso by contracting all edges in the separators $S(A_{i},B_{i})$. If all $(A_{i},B_{i})$ are ${}^{+}$, then the compressed torso and the expanded torso coincide. If $N$ is a nested set of mixed-separations of $G$, then the (compressed/expanded) torsos of $N$ are the (compressed/expanded) torsos of the splitting stars of $N$. We remark that if $N$ is the set of induced separations of a tree-decomposition $(T,{{\mathcal{V}}})$ of $G$, then the torsos of $N$ are precisely the torsos of $(T,{{\mathcal{V}}})$.
@@
2.2.4. Statement of the main theorem
Let $\sigma$ be a splitting star of a nested set $N$ of tri-separations of $G$. We say that a strong nontrivial tri-separation $(A,B)$ of $G$ interlaces $\sigma$ lightly if both $G[A\setminus B]$ and $G[B\setminus A]$ have at least two components. Otherwise $(A,B)$ interlaces $\sigma$ heavily. We stress that tri-separations that fail to be strong or nontrivial interlace $\sigma$ neither lightly nor heavily by definition.
@@ 2.2.4. Statement of the main theorem
Theorem 2.2.1.
@@
Figure 2.3. A graph with a thickened $K_{3,3}$ as torso; see Section 2.2.4
@@
For every integer $m\geqslant 0$, there exist $G$ and $\sigma$ as in the statement of Theorem 2.2.1 such that (i) holds and the compressed torso of $\sigma$ is a thickened $K_{3,m}$. Indeed, let $m$ be given. Let $X$ and $Y$ be disjoint vertex sets of size $m$ and three, respectively. We let $G$ be the graph obtained from the complete bipartite graph on $(X,Y)$ by disjointly adding four triangles $\Delta_{1},\ldots,\Delta_{4}$ and joining each triangle $\Delta_{i}$ to the three vertices in $Y$ by a matching of size three. Then $\sigma:=\{\,(\Delta_{i},V(G\setminus\Delta_{i})):i\in[4]\,\}$ is a splitting star of $N$. The splitting star $\sigma$ is lightly interlaced by the tri-separation $(\Delta_{1}\cup\Delta_{2}\cup Y,V(G\setminus(\Delta_{1}\cup\Delta_{2}))\,)$. The compressed torso of $\sigma$ is obtained from $G$ by contracting all edges in the matchings between $Y$ and the triangles $\Delta_{i}$, so it is a thickened $K_{3,m}$.
Figure 2.4. This graph is obtained from a $K_{n}$ with $n=100$ by first attaching the three vertices $v_{1}$, $v_{2}$ and $v_{3}$ of degree three as illustrated, and then deleting all edges with both ends in the neighbourhood of a $v_{i}$, except one edge which lies as in the figure.
@@
In (iii), we cannot replace ‘quasi 4-connected’ by ‘internally 4-connected’. Indeed, let $G$ be the graph depicted in Figure 2.4. The graph $G$ has precisely one strong nontrivial tri-separation (up to flipping sides), which has the form $(A,B)=(v_{1}+x+y,V(G)-v_{1})$ for $xy$ the unique edge with both ends in the neighbourhood of $v_{1}$. Hence $\{\,(A,B)\,\}$ is a splitting star of $N$. The compressed torso of $\sigma$ is obtained from $G$ by making the neighbourhood of the vertex $v_{1}$ complete and then deleting $v_{1}$. Thus this compressed torso has a nontrivial tri-separation, similar to the tri-separation $(A,B)$ with ‘$v_{2}$’ taking the role of ‘$v_{1}$’. Hence the compressed torso is quasi 4-connected but not internally 4-connected, compare Section 1.2.
Figure 2.5. An expanded torso that is not quasi 4-connected; red edges are missing
@@
In (iii), we cannot replace ‘compressed torso’ by ‘expanded torso’. Indeed, let $G$ be the graph obtained from $K_{10}$ by picking a triangle $\Delta$ within and attaching a new triangle $\Delta^{\prime}$ to $\Delta$ via a matching, and then deleting the edges of $\Delta$; see Figure 2.5. The matching is a 3-edge cut of $G$ and determines a splitting star $\{\,(\Delta^{\prime},G\setminus\Delta^{\prime})\,\}$ of $N$. Hence $G$ is an expanded torso of $N$, but the splitting star is not interlaced by a tri-separation and $G$ is not quasi 4-connected.
2.3. Proof of (i)
Our strategy to prove (i) is to construct for every tri-separation $(A,B)$ that is not half-connected a splitting star $\sigma$ interlaced by $(A,B)$. The construction is explicit and allows us to deduce that this splitting star has the structure for (i). Then we apply Section 2.2.2 to deduce that every splitting star interlaced by $(A,B)$ must be equal to $\sigma$, completing the proof. The details are as follows.
@@ 2.3. Proof of (i)
Suppose now that $G\setminus U$ has at least four components. By Section 1.5, for every 3-separation $(A_{K},B_{K})$ in this star the reduction $(A_{K}^{\prime},B_{K}^{\prime})$, which in particular is a tri-separation, satisfies $(A_{K}^{\prime},B_{K}^{\prime})\leqslant(A_{K},B_{K})$. The star of tri-separations induced by $U$ is the star that consists of all the reductions of the 3-separations in the star of 3-separations induced by $U$.
@@
Let $G$ be a 3-connected graph and $U\subseteq V(G)$ of size three such that $G\setminus U$ has at least three components. Then $U\subseteq A$ or $U\subseteq B$ for every nontrivial tri-separation of $G$.
@@
Let $U=:\{u,v,w\}$. Suppose for a contradiction that there is a nontrivial tri-separation $(A,B)$ of $G$ with $u\in A\setminus B$ and $v\in B\setminus A$. Since $G\setminus U$ has at least three components, and since every component $K$ has neighbourhood equal to $U$ by 3-connectivity, there exist three independent $u$$v$ paths $P_{K}$ in $G$ with all internal vertices included in $K$. Hence $S(A,B)$ must contain an edge or an internal vertex from each path $P_{K}$, and therefore $w$ cannot be contained in $S(A,B)$. So $w\in B\setminus A$, say. Hence $u$ is a cutvertex of $G[A]$. As $(A,B)$ is nontrivial, there is a cycle in $G[A]$. Let $K^{*}$ denote the component of $G\setminus U$ that contains a vertex of this cycle. The cycle has a vertex $x$, say, that is distinct from $u$ and not in $S(A,B)$ (since otherwise $S(A,B)$ would intersect $K^{*}$ in two vertices, a contradiction). But then the element of $S(A,B)$ on $P_{K^{*}}$ together with $u$ separate $x$ from the other components of $G\setminus U$, contradicting that $G$ is 3-connected. ∎
@@
By Section 2.3 we may assume that $U$ avoids $C\setminus D$, say. Hence $U\subseteq D$. If $S(C,D)=U$, then $(C,D)$ being half-connected implies that $(C,D)$ or $(D,C)$ is equal to a 3-separation $(A_{K},B_{K})$ for some component $K$ of $G\setminus U$, and we are done. So assume that $S(C,D)\neq U$. Let $K^{\prime}$ be an arbitrary component of $G[C\setminus D]$. Since $U$ is included in $D$, the component $K^{\prime}$ avoids $U$. So $K^{\prime}$ is included in a unique component $K$ of $G\setminus U$. We claim that $(C,D)\leqslant(A_{K},B_{K})$.
@@
The elements of $\sigma$ are totally-nested nontrivial tri-separations of $G$ by Section 1.5.
Suppose for a contradiction that $\sigma$ is not splitting. Then some totally-nested nontrivial tri-separation $(C,D)$ of $G$ interlaces $\sigma$ by Section 2.2.2. Since $(C,D)$ is totally nested, it is half-connected by Section 2.3. By Section 2.3 there is a component $K$ of $G\setminus U$ such that $(C,D)\leqslant(A_{K},B_{K})$, say. Let $(A_{K}^{\prime},B_{K}^{\prime})\in\sigma^{\prime}$ denote the reduction of $(A_{K},B_{K})$.
If $(A_{K}^{\prime},B_{K}^{\prime})$ is nontrivial, it lies in $\sigma$, so $(A_{K}^{\prime},B_{K}^{\prime})<(C,D)\leqslant(A_{K},B_{K})$ as $(C,D)$ interlaces $\sigma$. This contradicts Section 2.3. Hence $(A_{K}^{\prime},B_{K}^{\prime})$ is trivial; so $A_{K}^{\prime}$ and $B_{K}^{\prime}$ are the sides of an atomic 3-cut by Section 1.2. The only possibility here is that $|A_{K}^{\prime}|=1$, since $B_{K}^{\prime}$ includes $U$. But then $G[A_{K}]$ is a $K_{1,3}$, so $G[C]$ contains no cycle and $(C,D)$ is trivial, a contradiction. ∎
Proof of Theorem 2.2.1 (i)..
Let $\sigma$ be a splitting star of $N$ that is interlaced lightly by a tri-separation $(A,B)$ of $G$. Then $U=A\cap B$ is a 3-separator of $G$ such that $G\setminus U$ has at least four components. Let $\bar{\sigma}$ be the star of tri-separations induced by $U$, and note that $(A,B)$ interlaces $\bar{\sigma}$ as well. Let $\sigma^{\prime}$ consist of the nontrivial tri-separations in $\bar{\sigma}$. By Section 2.3, $\sigma^{\prime}$ is a splitting star of $N$. As $(A,B)$ interlaces the splitting stars $\sigma$ and $\sigma^{\prime}$ of $N$, these splitting stars need to be equal by Section 2.2.2.
If $G=K_{3,m}$ for some $m\geqslant 4$, then we are done. So we may assume that the graph $G[U]$ has an edge or $G\setminus U$ has a component of size at least two. Thus $\sigma^{\prime}$ is non-empty by Section 1.5. The compressed torso of $\sigma^{\prime}$ can be obtained from $G$ by first removing every component $K$ of $G\setminus U$ with $|K|\geqslant 2$, then removing every component $K$ of $G\setminus U$ with $|K|=1$ if $G[U]$ has an edge, and finally making $U$ into a clique (for the latter we need that $\sigma^{\prime}$ is nonempty). Hence the compressed torso of $\sigma=\sigma^{\prime}$ is a thickened $K_{3,m}$ with $m\geqslant 0$. ∎
2.4. Tools to prove (ii) and (iii)

### v3

@@ 2 Splitting stars
Let $(T,{{\mathcal{V}}})$ be a tree-decomposition of $G$, and let $S$ denote the set of induced separations of $(T,{{\mathcal{V}}})$. For every node $t\in T$, let $\sigma_{t}$ denote the star of separations induced by the edges of $T$ incident with $t$ and directed to $t$. The splitting stars of $S$ are precisely the stars $\sigma_{t}$ with $t\in T$.
@@
(1$\to$(2). Let $\sigma$ be a splitting star of $N$, and assume for a contradiction that $(C,D)\in N$ interlaces $\sigma$. Then there is $(A,B)\in\sigma$ such that $(C,D)\leqslant(A,B)$ or $(D,C)\leqslant(A,B)$. But we also have $(A,B)<(C,D)$ or $(A,B)<(D,C)$ since $(C,D)$ interlaces $\sigma$. In two cases we obtain immediate contradictions, and in the other two cases we obtain $C\subseteq D$ or $D\subseteq C$ which contradicts the definition of a separation.
(2$\to$(1). Assume that no element of $N$ interlaces $\sigma$; we show that $\sigma$ is a splitting star of $N$. Let $(C,D)\in N$, and assume for a contradiction that there is no $(A,B)\in\sigma$ such that $(C,D)\leqslant(A,B)$ or $(D,C)\leqslant(A,B)$. Then, since $N$ is nested, for every $(A,B)\in\sigma$ we have $(A,B)<(C,D)$ or $(A,B)<(D,C)$, so $(C,D)$ interlaces $\sigma$. ∎
@@
Assume for a contradiction that some mixed-separation $(A,B)$ of $G$ interlaces two distinct splitting stars $\sigma_{1}$ and $\sigma_{2}$ of $M$. By Section 2, there exist $(C_{1},D_{1})\in\sigma_{1}$ and $(C_{2},D_{2})\in\sigma_{2}$ such that $(D_{1},C_{1})\leqslant(C_{2},D_{2})$. Since $(A,B)$ interlaces $\sigma_{1}$ and $\sigma_{2}$, and since $(A,B)\not<(A,B)$ nor $(B,A)\not<(B,A)$, we either have
@@
3 Torsos
mixed-separation+ of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$ and no two edges in $E(A\setminus B,B\setminus A)$ share endvertices. We stress that we allow $A\setminus B$ and $B\setminus A$ to be empty. All the usual concepts for mixed-separations extend to mixed-separations+ in the obvious way.
@@ 3 Torsos
All nontrivial mixed 3-separations of a 3-connected graph are mixed 3-separations+ by Section 3. Pairs $(A,V(G))$ for $A\subseteq V(G)$ are separations+ but not separations.
Let $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ be a star of mixed-separations+ of a graph $G$, with leaf-sides $A_{i}$. The bag of $\sigma$ is the intersection $\bigcap_{i\in I}B_{i}$ of all non-leaf sides $B_{i}$. We follow the convention that the bag of the empty star is equal to the vertex-set of $G$.
If all $(A_{i},B_{i})$ are separations+, then the torso of $\sigma$ is the graph obtained from $G[\beta]$, where $\beta$ is the bag of $\sigma$, by making every separator $A_{i}\cap B_{i}$ into a clique (by adding all possible edges inside $A_{i}\cap B_{i}$ for all $i\in I$). In general, however, there are (at least) two ways how the notion of a torso can be generalised to stars of mixed-separations+. Here we present two ways, supported by Figure 15.
@@
The expanded torso of $\sigma$ is the graph that is obtained from $G$ as follows. We obtain $(A_{i}^{\prime},B_{i}^{\prime})$ from $(A_{i},B_{i})\in\sigma$ by letting $A_{i}^{\prime}:=A_{i}$ and we obtain $B_{i}^{\prime}$ from $B_{i}$ by adding all endvertices of edges in the separator of $(A_{i},B_{i})$. Then $(A_{i}^{\prime},B_{i}^{\prime})$ is a separation+ with the same order as $(A_{i},B_{i})$. We take the torso of the star $\{\,(A_{i}^{\prime},B_{i}^{\prime}):i\in I\,\}$ of separations+ as the expanded torso of $\sigma$.
Figure 15: Left: the star $\sigma=\{\,(A,B),(C,D)\,\}$. Middle: the expanded torso of $\sigma$. Right: the compressed torso of $\sigma$.
Note that the compressed torso can be obtained from the expanded torso by contracting all edges in the separators $S(A_{i},B_{i})$. If all $(A_{i},B_{i})$ are separations+, then the compressed torso and the expanded torso coincide. If $N$ is a nested set of mixed-separations of $G$, then the (compressed/expanded) torsos of $N$ are the (compressed/expanded) torsos of the splitting stars of $N$. We remark that if $N$ is the set of induced separations of a tree-decomposition $(T,{{\mathcal{V}}})$ of $G$, then the torsos of $N$ are precisely the torsos of $(T,{{\mathcal{V}}})$.
@@
4 Statement of the main theorem
Let $\sigma$ be a splitting star of a nested set $N$ of tri-separations of $G$. Let $(A,B)$ be a strong nontrivial tri-separation of $G$. We say that $(A,B)$ interlaces $\sigma$ lightly if $(A,B)$ interlaces $\sigma$ and both $G[A\setminus B]$ and $G[B\setminus A]$ have at least two components. We say that $(A,B)$ interlaces $\sigma$ heavily if $(A,B)$ interlaces $\sigma$ and $(A,B)$ is half-connected. So if $(A,B)$ interlaces $\sigma$, then it does so either lightly or heavily. We stress that tri-separations that fail to be strong or nontrivial interlace $\sigma$ neither lightly nor heavily by definition.
@@ 4 Statement of the main theorem
Recall that a graph $G$ is quasi 4-connected if $G$ is 3-connected, $G$ has $>4$ vertices, and every 3-separation of $G$ has a side of size at most four.
Theorem 2.1.
@@
Figure 16: A graph with a thickened $K_{3,3}$ as torso; see Section 4
@@
For every integer $m\geqslant 0$, there exist $G$ and $\sigma$ as in the statement of Theorem 2.1 such that (i) holds and the compressed torso of $\sigma$ is a thickened $K_{3,m}$. Indeed, let $m$ be given. Let $X$ and $Y$ be disjoint vertex sets of size $m$ and three, respectively. We let $G$ be the graph obtained from the complete bipartite graph on $(X,Y)$ by disjointly adding four triangles $\Delta_{1},\ldots,\Delta_{4}$ and joining each triangle $\Delta_{i}$ to the three vertices in $Y$ by a matching of size three. Then $\sigma:=\{\,(\Delta_{i},V(G\setminus\Delta_{i})):i\in[4]\,\}$ is a splitting star of $N$. The splitting star $\sigma$ is lightly interlaced by the tri-separation $(\Delta_{1}\cup\Delta_{2}\cup Y,V(G\setminus(\Delta_{1}\cup\Delta_{2}))\,)$. The compressed torso of $\sigma$ is obtained from $G$ by contracting all edges in the matchings between $Y$ and the triangles $\Delta_{i}$, so it is a thickened $K_{3,m}$.
Figure 17: This graph is obtained from a $K_{n}$ with $n=100$ by first attaching the three vertices $v_{1}$, $v_{2}$ and $v_{3}$ of degree three as illustrated, and then deleting all edges with both ends in the neighbourhood of a $v_{i}$, except one edge which lies as in the figure.
@@
In (iii), we cannot replace ‘quasi 4-connected’ by ‘internally 4-connected’. Indeed, let $G$ be the graph depicted in Figure 17. The graph $G$ has precisely one strong nontrivial tri-separation (up to flipping sides), which has the form $(A,B)=(v_{1}+x+y,V(G)-v_{1})$ for $xy$ the unique edge with both ends in the neighbourhood of $v_{1}$. Hence $\{\,(A,B)\,\}$ is a splitting star of $N$. The compressed torso of $\sigma$ is obtained from $G$ by making the neighbourhood of the vertex $v_{1}$ complete and then deleting $v_{1}$. Thus this compressed torso has a nontrivial tri-separation, similar to the tri-separation $(A,B)$ with ‘$v_{2}$’ taking the role of ‘$v_{1}$’. Hence the compressed torso is quasi 4-connected but not internally 4-connected, compare Section 3.
Figure 18: An expanded torso that is not quasi 4-connected; red edges are missing
@@
In (iii), we cannot replace ‘compressed torso’ by ‘expanded torso’. Indeed, let $G$ be the graph obtained from $K_{10}$ by picking a triangle $\Delta$ within and attaching a new triangle $\Delta^{\prime}$ to $\Delta$ via a matching, and then deleting the edges of $\Delta$; see Figure 18. The matching is a 3-edge cut of $G$ and determines a splitting star $\{\,(\Delta^{\prime},G\setminus\Delta^{\prime})\,\}$ of $N$. Hence $G$ is an expanded torso of $N$, but the splitting star is not interlaced by a tri-separation and $G$ is not quasi 4-connected.
3 Proof of (i)
Our strategy to prove (i) is to construct for every tri-separation $(A,B)$ that is not half-connected a splitting star $\sigma$ interlaced by $(A,B)$. The construction is explicit and allows us to deduce that this splitting star has the structure for (i). Then we apply Section 2 to deduce that every splitting star interlaced by $(A,B)$ must be equal to $\sigma$, completing the proof. The details are as follows.
@@ 3 Proof of (i)
Suppose now that $G\setminus U$ has at least four components. By Section 6, for every 3-separation $(A_{K},B_{K})$ in this star the reduction $(A_{K}^{\prime},B_{K}^{\prime})$, which in particular is a tri-separation, satisfies $(A_{K}^{\prime},B_{K}^{\prime})\leqslant(A_{K},B_{K})$. The star of tri-separations induced by $U$ is the star that consists of all the reductions of the 3-separations in the star of 3-separations induced by $U$.
@@
Let $G$ be a 3-connected graph and $U\subseteq V(G)$ of size three such that $G\setminus U$ has at least three components. Then $U\subseteq A$ or $U\subseteq B$ for every nontrivial tri-separation $(A,B)$ of $G$.
@@
Let $U=:\{u,v,w\}$. Suppose for a contradiction that there is a nontrivial tri-separation $(A,B)$ of $G$ with $u\in A\setminus B$ and $v\in B\setminus A$. Since $G\setminus U$ has at least three components, and since every component $K$ has neighbourhood equal to $U$ by 3-connectivity, there exist three independent $u$$v$ paths $P_{K}$ in $G$ with all internal vertices included in $K$. Hence $S(A,B)$ must contain an edge or an internal vertex from each path $P_{K}$. Therefore $w$ cannot be contained in $S(A,B)$ and $G\setminus U$ has exactly three components. So $w\in B\setminus A$, say. Hence $u$ is a cutvertex of $G[A]$. As $(A,B)$ is nontrivial, there is a cycle in $G[A]$. Let $K^{*}$ denote the component of $G\setminus U$ that contains a vertex of this cycle. The cycle has a vertex $x$, say, that is distinct from $u$ and not in $S(A,B)$ (since otherwise $S(A,B)$ would intersect $K^{*}$ in two vertices, a contradiction). But then the element of $S(A,B)$ on $P_{K^{*}}$ together with $u$ separate $x$ from the other components of $G\setminus U$, contradicting that $G$ is 3-connected. ∎
@@
By Section 3 we may assume that $U$ avoids $C\setminus D$, say. Hence $U\subseteq D$. If $S(C,D)=U$, then $(C,D)$ being half-connected implies that $(C,D)$ or $(D,C)$ is equal to a 3-separation $(A_{K},B_{K})$ for some component $K$ of $G\setminus U$, and we are done. So assume that $S(C,D)\neq U$. Let $K^{\prime}$ be an arbitrary component of $G[C\setminus D]$. Since $U$ is included in $D$, the component $K^{\prime}$ avoids $U$. So $K^{\prime}$ is included in a unique component $K$ of $G\setminus U$. We claim that $(C,D)\leqslant(A_{K},B_{K})$.
@@
The elements of $\sigma$ are totally-nested nontrivial tri-separations of $G$ by Section 6.
Suppose for a contradiction that $\sigma$ is not splitting. Then some totally-nested nontrivial tri-separation $(C,D)$ of $G$ interlaces $\sigma$ by Section 2. Since $(C,D)$ is totally nested, it is half-connected by Section 3. By Section 3 there is a component $K$ of $G\setminus U$ such that $(C,D)\leqslant(A_{K},B_{K})$, say. Let $(A_{K}^{\prime},B_{K}^{\prime})\in\sigma^{\prime}$ denote the reduction of $(A_{K},B_{K})$.
If $(A_{K}^{\prime},B_{K}^{\prime})$ is nontrivial, it lies in $\sigma$, so $(A_{K}^{\prime},B_{K}^{\prime})<(C,D)\leqslant(A_{K},B_{K})$ as $(C,D)$ interlaces $\sigma$. This contradicts Section 3. Hence $(A_{K}^{\prime},B_{K}^{\prime})$ is trivial; so $A_{K}^{\prime}$ and $B_{K}^{\prime}$ are the sides of an atomic 3-cut by Section 3. The only possibility here is that $|A_{K}^{\prime}|=1$, since $B_{K}^{\prime}$ includes $U$. But then $G[A_{K}]$ is a $K_{1,3}$, so $G[C]$ contains no cycle and $(C,D)$ is trivial, a contradiction. ∎
Proof of Theorem 2.1 (i)..
Let $\sigma$ be a splitting star of $N$ that is interlaced lightly by a tri-separation $(A,B)$ of $G$. Then $U=A\cap B$ is a 3-separator of $G$ such that $G\setminus U$ has at least four components. Let $\bar{\sigma}$ be the star of tri-separations induced by $U$, and note that $(A,B)$ interlaces $\bar{\sigma}$ as well. Let $\sigma^{\prime}$ consist of the nontrivial tri-separations in $\bar{\sigma}$. By Section 3, $\sigma^{\prime}$ is a splitting star of $N$. As $(A,B)$ interlaces the splitting stars $\sigma$ and $\sigma^{\prime}$ of $N$, these splitting stars need to be equal by Section 2.
If $G=K_{3,m}$ for some $m\geqslant 4$, then we are done. So we may assume that the graph $G[U]$ has an edge or $G\setminus U$ has a component of size at least two. Thus $\sigma^{\prime}$ is non-empty by Section 6. The compressed torso of $\sigma^{\prime}$ can be obtained from $G$ by first removing every component $K$ of $G\setminus U$ with $|K|\geqslant 2$, then removing every component $K$ of $G\setminus U$ with $|K|=1$ if $G[U]$ has an edge, and finally making $U$ into a clique (for the latter we need that $\sigma^{\prime}$ is nonempty). Hence the compressed torso of $\sigma=\sigma^{\prime}$ is a thickened $K_{3,m}$ with $m\geqslant 0$. ∎
4 Tools to prove (ii) and (iii)